# MoESD: Mixture of Experts Stable Diffusion to Mitigate Gender Bias

### v1

@@ A.5 Parameter $\lambda$ for Bias-Identification
A.6 Defect of Bias Prompts
@@
Figure 7: Latent Representation of Prompt Embedding.
So Chuang et al. (2023) exhibits a defect that we have observed: as the parameter $\lambda$ increases, the original characteristics are lost and male characteristics become more and more prominent. We present a different set of images, where each set of images are generated with the same prompt and seeds but with different $\lambda$ to illustrate the “male guidance”.
Figure 8: Defect of Chuang et al. (2023). The male characteristics become more and more prominent, and in some cases, the gender even switches.
A.7 Occupation list
We use the occupation list from Fair Diffusion Friedrich et al. (2023) instead of Chuang et al. (2023), as we found that the Fair Diffusion one contains more occupations and includes more challenging ones. All occupations are displayed in the Table 6 below, where we specifically label the right and wrong occupation predictions for our Bias-Identification method.
@@ A.7 Occupation list
A.8 BLIP2 VQA for Fairness Evaluation
@@
A.9 Failure cases
A.9.1 Right and Wrong Case for Bias Identification
@@
A.9.2 Failure Generation for our Method
As mentioned above, some of our generated images blur the faces. This could be further improved through fine-tuning.
@@ A.9.2 Failure Generation for our Method
A.10 Human Evaluation Details

### v2

@@ Appendix E Parameter $\lambda$ for Bias-Identification
Appendix F Defect of DebiasVL
@@
Figure 7: Latent Representation of Prompt Embedding
So DebiasVL exhibits a defect that we have observed: as the parameter $\lambda$ increases, the original characteristics are lost and male characteristics become more and more prominent. We present a different set of images, where each set of images is generated with the same prompt and seeds but with different $\lambda$ to illustrate the “male guidance”, shown in Figure 8.
Figure 8: Defect of DebiasVL. The male characteristics become more and more prominent, and in some cases, the gender even switches.
Appendix G Occupation list
We use the occupation list from Fair Diffusion instead of DebiasVL, as we found that the Fair Diffusion one contains more occupations and includes more challenging ones. All occupations are displayed in the Table 6 below, where we specifically label the right and wrong occupation predictions for our Bias-Identification method.
@@ Appendix G Occupation list
Appendix H BLIP2 VQA for Fairness Evaluation
Followed by Esposito et al. (2023), we use BLIP2 VQA for gender mitigation evaluation and skin tone supplemental experiment.
H.1 Gender
@@
H.2 Skin Tone
For each picture, we use the following question to conduct the VQA evaluation and count the number of skin tone in 100 pictures for each occupation.
Question: “ Answer the following question with ‘light’ or ‘medium’ or ‘black’ (dark skin) or ‘people not present’ only. What is the skin tone of this person?”
However, the model is not willing to classify skin tone as ‘medium’ or ‘people not present’; it only categorizes them as either ‘light’ or ‘dark’.
Appendix I Failure cases
I.1 Right and Wrong Case for Bias Identification
@@
I.2 Blurring Generation in Our Method
Some of our generated images exhibit blurred faces. We present our human evaluation results for aesthetics, showing that our method was rated the same or better than the vanilla method only in 54.6% of the cases. This reflects the fact that our images are sometimes blurry, a very noticeable artifact that often leads our human evaluators to (rightly) prefer vanilla images. This lower quality may be due to the limited size and synthetic nature of the training data, as well as the small number of training parameters. This indicates that more work is needed to maintain the quality of the generated images while mitigating biases. Improvements could be achieved through more rigorous fine-tuning and the use of high-quality datasets.
@@ I.2 Blurring Generation in Our Method
Appendix J Human Evaluation Details
@@
Appendix K Dataset and Computing Infrastructure Explanation
Our dataset uses the prompt list from Fair Diffusion, which includes a list of 153 occupations as shown in Appendix G. The experiment was conducted on NVIDIA RTX A6000 (48 GB) and Tesla V100-PCIE (32 GB) GPUs, running on Linux kernel version 6.5.0-41-generic.
Appendix L Randomness and Number of Algorithm Runs
We set the random seeds for initial noise so that our results, as well as the DebiasVL results and vanilla results, remain consistent regardless of the number of runs. For Ethical Intervention and Fairness Finetuning, we perform a single run.
Appendix M Limitations
Although our method has made significant strides in mitigating bias while maintaining aesthetic quality, we must acknowledge the limitations of our work.
First, we only address the case of binary gender with two bias experts. Considering sexual minorities in various contexts is a much more complex task and might need more precise bias identification methods and additional experts. We hope to explore this further in future work.
Second, our zero-shot and unsupervised prompt bias identification and hyperparameter selection are based on the statistical counts from Fair Diffusion (Friedrich et al. 2023), which may not be entirely accurate. Moreover, we only achieve 79% accuracy in identifying bias from the prompt, which is not perfect.
Third, although we achieve good performance in aesthetics, our method does lose some fine-grained details on faces due to the quality of the fine-tuning image dataset. However, fortunately, our experts can be replaced with fine-tuned ones by users themselves, which can be further improved through specialized fine-tuning.
Appendix N Ethical Considerations
However, by introducing the BiAs approach to mitigate bias, there is a risk that people might misuse the weights to generate more biased content. Moreover, in our work, we only address the case of binary gender. We did not consider sexual minorities in various contexts, which is a much more complex task. We hope to explore this further in future work.